Lance Whittingham IV specializes in buying deep discount bonds. These represent bonds that are trading at well below par value. He has his eye on a bond issued by the Leisure Time Corporation. The $1,000 par value bond pays 7 percent annual interest and has 16 years remaining to maturity. The current yield to maturity on similar bonds is 15 percent.
  
a. What is the current price of the bonds? Use Appendix B and Appendix D for an approximate answer but calculate your final answer using the formula and financial calculator methods. (Do not round intermediate calculations. Round your final answer to 2 decimal places. Assume interest payments are annual.)
  


   
b. By what percent will the price of the bonds increase between now and maturity? (Do not round intermediate calculations. Input your answer as a percent rounded to 2 decimal places.)
  


  

eBook & Resources

eBook: Valuation of Bonds
Answer #1

Given,

FV = 1000

PMT = 0.07*1000 = 70

N = 16

I = 0.15

Input these values in financial calculator, compute, CPT PV = 523.66

Using formula:

Rate 0.1500
Year(n) Cashflow (A) Discount rate = 1/(1+r)^(n) Present Value = A*discount rate
1 70 0.86956522 60.86956522
2 70 0.75614367 52.93005671
3 70 0.65751623 46.02613627
4 70 0.57175325 40.02272719
5 70 0.49717674 34.80237147
6 70 0.4323276 30.26293171
7 70 0.37593704 26.31559279
8 70 0.32690177 22.88312417
9 70 0.28426241 19.89836884
10 70 0.24718471 17.30292943
11 70 0.21494322 15.04602559
12 70 0.18690715 13.08350051
13 70 0.16252796 11.37695697
14 70 0.14132866 9.893006059
15 70 0.12289449 8.602613964
16 1070 0.10686477 114.3453036
NPV 523.6612105

b.

At maturity, bond price = par value

Bond price = 1000

Increase = 1000 - 523.66 = 476.34

% increase = 476.34/523.66 = 0.9096 = 90.96%